# Building Instruction-Tuning Datasets from Human-Written Instructions with Open-Weight Large Language Models

### v1

@@ 2 Synthesized Data Generation
We collected instructions =$\mathcal{I}=\{I_{1},I_{2},\cdots,I_{n}\}$ from LMSYS-Chat-1M (Zheng et al., 2024) [^2] https://huggingface.co/datasets/lmsys/lmsys-chat-1m, a dataset of one million real-world conversation records between human and LLMs. The dataset was selected as our foundation as it covers a wide range of topics, with a permissive license for both academic and commercial use. Notably, some of the conversations in LMSYS-Chat-1M are flagged as toxic; these conversations were excluded as our focus here is to improve LLMs’ instruction-following abilities, i.e., how accurately LLMs can fulfill user requests. After filtering out toxic conversations, we obtained 732,392 instances, from which the first user utterance is selected as the instruction. Further removal of duplications left us with 453,889 instructions directly adopted for synthesizing responses.
@@
Statistics of our datasets, namely {Llama-3.1|Gemma-2}-Proposed-En, are reported in Table 1. A dataset $D$ synthesized by an LLM $\hat{\pi}$ enables models to imitate how $\hat{\pi}$ responds to human-written instructions. We hereafter refer to the LLM utilized for response generation as a teacher model and the LLM fine-tuned on the dataset as a student model.
@@ 3 Experiments
Llama-3.1-405B-Instruct 8.33±.$\pm.08$ 8.39±.$\pm.08$ 8.27±.$\pm.12$
Gemma-2-27B-IT 7.89±.$\pm.05$ 8.19±.$\pm.06$ 7.54±.$\pm.10$
Llama-3.1-8B LMSYS-Chat-1M-Random 4.10±.$\pm.08$ 4.77±.$\pm.15$ 3.36±.$\pm.14$
LMSYS-Chat-1M-Best 4.15±.$\pm.08$ 4.89±.$\pm.05$ 3.36±.$\pm.13$
Ultra-Chat 5.13±.$\pm.11$ 5.76±.$\pm.05$ 4.43±.$\pm.18$
Magpie-Ultra-v0.1 5.61±.$\pm.11$ 6.32±.$\pm.06$ 4.84±.$\pm.27$
Magpie-Gemma2-Pro-Filtered 5.82±.$\pm.12$ 6.30±.$\pm.07$ 5.24±.$\pm.21$
Magpie-Llama-3.1-Pro-MT-Filtered 6.40±.$\pm.08$ 6.86±.$\pm.11$ 5.90±.$\pm.16$
@@
Proposed-Gemma-2-En 6.40±.$\pm.10$ 6.58±.$\pm.06$ 6.21±.$\pm.17$
Gemma-2-9B Magpie-Llama-3.1-Pro-MT-Filtered 6.47±.$\pm.08$ 7.23±.$\pm.03$ 5.65±.$\pm.16$
@@
Proposed-Gemma-2-En 6.58±.$\pm.06$ 7.11±.$\pm.03$ 5.97±.$\pm.12$
Qwen-2.5-7B Magpie-Llama-3.1-Pro-MT-Filtered 5.61±.$\pm.05$ 6.33±.$\pm.07$ 4.82±.$\pm.07$
@@
Proposed-Gemma-2-En 6.92±.$\pm.06$ 7.59±.$\pm.07$ 6.17±.$\pm.13$
OLMo-2-1124-13B Magpie-Llama-3.1-Pro-MT-Filtered 6.85±.$\pm.10$ 7.16±.$\pm.02$ 6.50±.$\pm.21$
Tülu 3 5.92±.$\pm.16$ 6.41±.$\pm.13$ 5.38±.$\pm.22$
@@ 3.1 Configurations
We conducted full-parameter Supervised Fine-Tuning (SFT) for every instruction-tuning experiment reported in this paper. Specifically, all model parameters were updated to minimize the log probability of response $R_{k}$ given instruction $I_{k}$. The training process utilizes Huggingface’s transformers library (Wolf et al., 2020), and distributed across 4 or 8 NVIDIA H100 SXM5 GPUs using DeepSpeed ZeRO (Rajbhandari et al., 2020). All experiments were finished in 24 hours except those with Gemma-2-9B, which took 40 hours at maximum. AdamW (Loshchilov & Hutter, 2019) was adopted as the optimizer, with $\beta_{1}=0.9$ and $\beta_{2}=0.95$. The learning rate was adjusted following a cosine annealing schedule, which linearly increased to the peak at $2.5\times 10^{-5}$ in the first 10% number of steps, then smoothly decreased to $2.5\times 10^{-6}$. Each training process ran for 2 epochs with batch size $512$ across all experiments.
@@ 3.2 Main Results
The results are presented in Table 2. We detail the results of Llama-3.1-8B tuned on each baseline, strong competitor, and our dataset. For other student models, only the strongest competitor, i.e., Magpie-Llama-3.1-Pro-MT-Filtered is included due to page constraints.
@@ 4.1 Configurations
We adopt Japanese MT-Bench, a variant of MT-Bench that reflects Japan’s language, education, society, and culture [^6] https://wandb.ai/wandb-japan/llm-leaderboard/reports/Nejumi-LLM-Neo--Vmlldzo2MTkyMTU0. Japanese is one of the few languages which has reliable evaluations baselines, which motivated our choice.
@@ 4.2 Results
  Model SFT Data MT-Bench
@@
Gemma-2-27B-IT 7.10
Llama-3.1-405B-Instruct 6.79
Llama-3.1-8B official 4.65
Ja-Self-Inst 3.85
Proposed-Gemma-2-Ja 5.64
Proposed-Llama-3.1-Ja 5.05
Gemma-2-9B official 6.75
Ja-Self-Inst 3.98
Proposed-Gemma-2-Ja 6.49
Proposed-Llama-3.1-Ja 5.14
Qwen-2.5-7B official 6.13
Ja-Self-Inst 4.24
Proposed-Gemma-2-Ja 6.39
Proposed-Llama-3.1-Ja 5.40
Llama-3.1-Swallow-8B official 5.33
Ja-Self-Inst 4.53
Proposed-Gemma-2-Ja 6.23
Proposed-Llama-3.1-Ja 5.18
llm-jp-3-13b official 5.18
Ja-Self-Inst 3.71
Proposed-Gemma-2-Ja 5.38
Proposed-Llama-3.1-Ja 4.74
Table 4: Model performance evaluated on Japanese MT-Bench. 4 out of 5 LLMs show better performance when fine-tuned on our dataset than their official post-tuned release.
@@
Next, we observe the dataset synthesized with Gemma-2-27B-IT functions much better than that synthesized with Llama-3.1-405B-Instruct. While Gemma-2-27B-IT merely outperforms Llama-3.1-405B-instruct by 0.32 points, the gap between student models fine-tuned on synthesized datasets is larger. This is counterintuitive, as we have witnessed the performance of the student model being proportional to that of the teacher model in English. This indicates that the effectiveness of model imitation for non-English languages does not necessarily align with the findings in English.
@@ 5 Analysis
Our datasets enable cross-lingual analyses of instruction tuning. Here, we analyze how language capabilities are enhanced after instruction-tuning (IT) and distinguish them from those enhanced after continuous pre-training (CPT).
Purpose.
Pretraining LLMs per language from scratch demands extensive data and computation resources (Zeng et al., 2023; Aizawa et al., 2024); to avoid this, research has been conducted to adapt LLMs to underrepresented languages via IT (Chen et al., 2024; Shaham et al., 2024; Zhang et al., 2024). Concurrent to this work, Wang et al. (2025) have reported that incorporating CPT before IT better improves the capability in the target language. However, the role of CPT and IT in enhancing language capabilities remains unclear. We decouple the language capabilities into two categories, namely language understanding and language utilization, and investigate how CPT and IT contribute to acquiring these capabilities.
@@
(a) Model performance evaluated on Japanese academic tasks and MT-Bench. Scores of MT-Bench are normalized to 0–1. IT(JA) enhances performance on MT-Bench but does not help academic tasks.
(b) Model performance on each sub-category of Japanese MT-Bench. IT helps writing, stem, and roleplay, while CPT contributes to humanities.
@@
On academic tasks, we observe significant performance improvement in models who underwent CPT. In contrast, models without CPT show only subtle changes in performance, even after IT with the Japanese dataset. These results suggest that CPT is crucial in equipping LLMs with fundamental language understanding abilities.
@@ 7 Conclusion
While not investigated within the scope of this work, we plan to further evolve the utility of our datasets by adding support for safety and multi-turn conversations as future work.
@@
This paper focuses more on improving the usefulness than the safety of LLMs. As a result, although we filtered out instructions flagged as toxic in LMSYS-Chat-1M, we acknowledge that some toxic contents still remain in our datasets. LLMs trained with our datasets should undergo further fine-tuning or reinforcement learning to ensure their safety and fairness.
Derived from LMSYS-Chat-1M (Zheng et al., 2024), all instructions in our datasets have tokens that could reveal Personally Identifiable Information (PII) masked out. For LLM-generated responses, we leave them as-is, as the synthetic data generation process does not introduce any additional information about real individuals.
@@ Acknowledgments
The research and development of the large language model Swallow has been supported by the AIST Project "Research and Development on Generative AI Foundation Models in the Physical Domain," the "Core Integrated Technology Development for Next-Generation Artificial Intelligence and Robotics" project by the New Energy and Industrial Technology Development Organization (NEDO) (JPNP18002), specifically focusing on "Development of AI Application Technology for Decision Support in Design Risk Assessment Based on Expert Perspectives." It is also supported by a project from the Ministry of Education, Culture, Sports, Science, and Technology (MEXT) aimed at "establishment of research and development centers to ensure the transparency and reliability of generative AI models", along with other contributions. This research was conducted using the ABCI large-scale generative AI research and development support program and the TSUBAME4.0 supercomputer at Institute of Science Tokyo. We thank Sangwhan Moon for his valuable suggestions on improving the manuscript.
Appendix A License Information of our datasets
@@
Appendix B Task-Wise Effectiveness
@@
MT-Bench includes instructions across eight categories. Here, we examine how our collected datasets enhance performance in each category. Evaluation results for LLMs tuned on Proposed-Llama-3.1-En and Proposed-Gemma-2-Ja are shown in Figure 3(a) and 3(b), respectively. The performance of teacher models, i.e., LLMs utilized for generating responses, is also included to see how the student models inherited the capabilities.
@@
On the other hand, in the math category for both languages, we observe that Qwen-2.5 leads in performance. The base models in the Qwen series are known for their strength in math tasks. Similarly, in the Japanese humanities category (Figure 3(b)), Llama-3.1-Swallow outperforms Llama-3.1 and demonstrates the best performance among all student models. Llama-3.1-Swallow, which was continuously pre-trained with Japanese corpus on Llama-3.1, gains a deeper understanding of Japanese local culture than Llama-3.1 Okazaki et al. (2024); Fujii et al. (2024). These results suggest that the inherent advantages in reasoning abilities and knowledge from base models persist even after instruction-tuning.
@@
Appendix C Instance-Wise Effectiveness: Japanese

### v2

@@ 2 Synthesized Data Generation
We collected instructions $\mathcal{I}=\{I_{1},I_{2},\cdots,I_{n}\}$ from LMSYS-Chat-1M (Zheng et al., 2024) [^2] https://huggingface.co/datasets/lmsys/lmsys-chat-1m, a dataset of one million real-world conversation records between human and LLMs. The dataset was selected as our foundation as it covers a wide range of topics, with a permissive license for both academic and commercial use. Notably, some of the conversations in LMSYS-Chat-1M are flagged as toxic; these conversations were excluded as our focus here is to improve LLMs’ instruction-following abilities, i.e., how accurately LLMs can fulfill user requests. After filtering out toxic conversations, we obtained 732,392 instances, from which the first user utterance is selected as the instruction. Further removal of duplications left us with 453,889 instructions directly adopted for synthesizing responses.
@@
Statistics of our datasets, {Llama-3.1|Gemma-2}-Proposed-En, are reported in Table 1. A dataset $D$ synthesized by an LLM $\hat{\pi}$ enables models to imitate how $\hat{\pi}$ responds to human-written instructions. We hereafter refer to the LLM utilized for response generation as a teacher model and the LLM fine-tuned on the dataset as a student model.
@@ 3.1 Configurations
We conducted full-parameter Supervised Fine-Tuning (SFT) for every instruction-tuning experiment reported in this paper. Specifically, all model parameters were updated to maximize the log probability of response $R_{k}$ given instruction $I_{k}$. The training process utilizes Huggingface’s transformers library (Wolf et al., 2020), and distributed across 4 or 8 NVIDIA H100 SXM5 GPUs using DeepSpeed ZeRO (Rajbhandari et al., 2020). All experiments were finished in 24 hours except those with Gemma-2-9B, which took 40 hours at maximum. AdamW (Loshchilov & Hutter, 2019) was adopted as the optimizer, with $\beta_{1}=0.9$ and $\beta_{2}=0.95$. The learning rate was adjusted following a cosine annealing schedule, which linearly increased to the peak at $2.5\times 10^{-5}$ in the first 10% number of steps, then smoothly decreased to $2.5\times 10^{-6}$. Each training process ran for 2 epochs with batch size $512$ across all experiments.
@@ 3.2 Main Results
The results are presented in Table 2. We detail the results of Llama-3.1-8B tuned on each baseline, strong competitor, and our dataset. For other student models, only the strongest competitor, i.e., Magpie-Llama-3.1-Pro-MT-Filtered is included.
@@ 4.1 Configurations
We adopt Japanese MT-Bench, a variant of MT-Bench that reflects Japan’s language, education, society, and culture [^6] https://wandb.ai/wandb-japan/llm-leaderboard/reports/Nejumi-LLM-Neo--Vmlldzo2MTkyMTU0. Japanese is one of few languages with reliable evaluation baselines, which motivated our choice. In addition, we measure the Japanese character ratio (Ja. char. ratio in Table 4) in the generated outputs to assess the input-output language consistency. A low Japanese character ratio indicates that the model generates output in off-target languages (Sennrich et al., 2024). During pilot experiments, we empirically found that a Japanese character ratio below 0.60 indicates a noticeable presence of content written in off-target languages.
@@ 4.2 Results
  Model SFT Data MT-Bench Ja. char. ratio
@@
[…]
Table 4: Model performance evaluated on Japanese MT-Bench. 4 out of 5 LLMs show better performance when fine-tuned on our dataset than their official post-trained release.
@@
Next, we observe that the dataset synthesized with Gemma-2-27B-IT functions much better than that synthesized with Llama-3.1-405B-Instruct. While Gemma-2-27B-IT merely outperforms Llama-3.1-405B-instruct by 0.32 points, the gap between student models fine-tuned on the corresponding synthesized datasets is larger. This is counterintuitive, as we have witnessed the performance of the student model being proportional to that of the teacher model in English. This indicates that the effectiveness of model imitation for non-English languages does not necessarily align with the findings in English.
In terms of the Japanese character ratio, models trained on our datasets consistently maintain a safe level above 0.60, matching that of their teacher models. Two native-level Japanese speakers with PhD degrees reviewed the responses generated by the teacher models and found off-target language in fewer than 0.5% of cases. Given that models fine-tuned on our datasets exhibit a similar Japanese character ratio as their corresponding teacher models, we conclude that the risk of these models producing responses in off-target languages is low.
@@ 5 Analysis
Figure 2: Counts of top-10 most frequent topics in Magpie-Llama-3.1-Pro-MT-Filtered (Xu et al., 2025) and Proposed-Llama-3.1-En. Topics in Magpie are skewed to mathematics, while those in the proposed dataset are more balanced.
5.1 Quantitative Analysis
We have shown that human-written instructions in our proposed dataset are superior to the LLM-generated instructions in Magpie (Xu et al., 2025). To better understand their difference, we dive into the instructions in Magpie-Llama-3.1-Pro-MT-Filtered and Proposed-Llama-3.1-En. We use InsTag (Lu et al., 2024) to annotate each user instruction with its semantic intent, referred to as the topic, to examine differences between Magpie and our dataset.
Figure 2 presents the top-10 most frequent topics and their counts. In Magpie, we observe a highly skewed distribution: the most frequent topic, namely mathematics, appears 1,520 times—nearly three times more than the second most frequent topic (530 times). In contrast, our proposed dataset exhibits a more uniform distribution across the top-10 topics, covering diverse categories such as article writing, grammar checking, and general reasoning. This analysis reveals that Magpie’s instruction set is heavily biased toward mathematical topics, whereas our dataset provides broader and more balanced topic coverage. Such diversity likely contributes to the superior performance of models fine-tuned on our data. We also observe a correlation between the category-wise performance of fine-tuned models and the topic distributions quantified in this section. Further details are provided in Appendix E.
5.2 Cross-Lingual Analysis
Our datasets enable cross-lingual analyses of instruction tuning. Here, we analyze how language capabilities are enhanced after instruction-tuning (IT) and distinguish them from those enhanced after continuous pre-training (CPT). Pretraining LLMs per language from scratch demands extensive data and computation resources (Zeng et al., 2023; Aizawa et al., 2024); to avoid this, research has been conducted to adapt LLMs to underrepresented languages via IT (Chen et al., 2024; Shaham et al., 2024; Zhang et al., 2024). Concurrent to this work, Wang et al. (2025) have reported that incorporating CPT before IT better improves the capability in the target language. However, the role of CPT and IT in enhancing language capabilities remains unclear. We decouple the language capabilities into two categories, namely language understanding and language utilization, and investigate how CPT and IT contribute to acquiring these capabilities.
@@ 5.2 Cross-Lingual Analysis
(a) Scores of MT-Bench are normalized to 0–1. IT(JA) enhances performance on MT-Bench but does not help academic tasks.
(b) IT helps writing, stem, and roleplay, while CPT contributes to humanities.
Figure 3: Model performance on each sub-category of Japanese MT-Bench.
@@
On academic tasks, we observe significant performance improvement in models that underwent CPT. In contrast, models without CPT show only subtle changes in performance, even after IT with the Japanese dataset. These results suggest that CPT is crucial in equipping LLMs with fundamental language understanding abilities.
@@ 7 Conclusion
While not explored in this study, we plan to enhance the utility of our datasets by incorporating support for safety and multi-turn conversations. Another future direction is to construct and evaluate the utility of synthesized datasets with instructions derived from human-written instructions that have been empirically validated as effective.
@@
This paper focuses more on improving the usefulness than the safety of LLMs. As a result, although we filtered out instructions flagged as toxic in LMSYS-Chat-1M, we acknowledge that some toxic contents still remain in our datasets. LLMs trained with our datasets should undergo further fine-tuning or reinforcement learning to ensure their safety and fairness. Derived from LMSYS-Chat-1M (Zheng et al., 2024), all instructions in our datasets have tokens that could reveal Personally Identifiable Information (PII) masked out. For LLM-generated responses, we leave them as-is, as the synthetic data generation process does not introduce any additional information about real individuals.
@@ Acknowledgments
This work was supported by JSPS KAKENHI Grant Number 25H01137. This work was also supported by JST K Program Japan Grant Number JPMJKP24C3. The research and development of the large language model Swallow has been supported by the AIST Project "Research and Development on Generative AI Foundation Models in the Physical Domain". It was also supported by a project from the Ministry of Education, Culture, Sports, Science, and Technology (MEXT) aimed at "establishment of research and development centers to ensure the transparency and reliability of generative AI models", along with other contributions. This research was conducted using the ABCI large-scale generative AI research and development support program, and the TSUBAME 4.0 supercomputer at Institute of Science Tokyo. We thank Sangwhan Moon for his valuable suggestions on improving the manuscript.
Appendix A License Information
@@
Appendix B Additional Experiments
This section presents additional experiments not included in Table 2, as shown in Table 5. Results for LMSYS-Chat-1M-Random, LMSYS-Chat-1M-Best (Zheng et al., 2024), and Ultra-Chat (Ding et al., 2023) are omitted, as they serve as a weak baseline intended for sanity checks; Models fine-tuned on these datasets significantly underperform those fine-tuned on Magpie (Xu et al., 2025) and our proposed datasets.
Table 5 shows trends consistent with Table 2, confirming the superiority of our proposed datasets over Magpie.
  MT-Bench
Base Model SFT Dataset Average 1st Turn 2nd Turn
      Reference: performance of teacher models
[…]
Table 5: Model performance evaluated on MT-Bench. $\dagger$ represents the official recipe utilized in the supervised fine-tuning of OLMo-2-1124-13B, with evaluation results of allenai/OLMo-2-1124-13B-SFT reported. This table is complementary to Table 2.
Appendix C Category-Wise Effectiveness
@@
MT-Bench includes instructions across eight categories. Here, we examine how our collected datasets enhance performance in each category. Evaluation results for LLMs tuned on Proposed-Llama-3.1-En and Proposed-Gemma-2-Ja are shown in Figure 4(a) and 4(b), respectively. The performance of teacher models, i.e., LLMs utilized for generating responses, is also included to see how the student models inherited the capabilities.
@@
On the other hand, in the math category for both languages, we observe that Qwen-2.5 leads in performance. The base models in the Qwen series are known for their strength in math tasks. Similarly, in the Japanese humanities category (Figure 4(b)), Llama-3.1-Swallow outperforms Llama-3.1 and demonstrates the best performance among all student models. Llama-3.1-Swallow, which was continuously pre-trained with Japanese corpus on Llama-3.1, gains a deeper understanding of Japanese local culture than Llama-3.1 (Okazaki et al., 2024; Fujii et al., 2024). These results suggest that the inherent advantages in reasoning abilities and knowledge from base models persist even after instruction-tuning.
@@
Appendix D Instance-Wise Effectiveness: Japanese
@@ Results.
Appendix E Category-Wise Quantitive Analysis
Figure 5: Category-wise scores of models fine-tuned on Magpie-Llama-3.1-Pro-MT-Filtered and Proposed-Llama-3.1-En. The performance gap is larger in categories underrepresented in Magpie’s topic distribution.
Corresponding to Section 5.1, here we report the category-wise performance of Llama-3.1-8B trained on Magpie-Llama-3.1-Pro-MT-Filered and Proposed-Llama-3.1-En in Figure 5.
As shown in the plot, models trained on our dataset outperform those trained on Magpie in the Humanities and Roleplay categories, while the advantage diminishes in Reasoning and Math. This trend may reflect the topic distributions discussed in Section 5.1, where our dataset is more balanced but Magpie is heavily skewed toward mathematics. The skew likely boosts performance in Reasoning and Math, consistent with recent findings on the close relationship between mathematical and reasoning abilities (DeepSeek-AI, 2025).